Construct 17−−√ on the number line.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Hint: We know that 16−−√=4 and 17−−√=42+12−−−−−−√. Squaring both sides of the expression and constructing it on a number line. Follow the steps to represent the number on a number line and get the required figure. Complete step by step answer: The number system represents the numbers on a number line. Number system is a very useful and important concept in mathematics to represent the numbers on a number line. All types of numbers like natural numbers, whole numbers, rational numbers, Integers are represented on a number line. All the numbers natural numbers, whole numbers, rational numbers, integers are collectively called as real numbers A number line is a line that serves as an abstraction of real numbers. Every point on a number line is assumed as a real number. The numbers on a number line are placed at equal intervals. According to the question it is given that the number is 17−−√. Here, 17−−√can be written as, ⇒17−−√=42+12−−−−−−√ ⇒17−−√=16+1−−−−−√ Squaring on both sides, ⇒(17−−√)2=42+12 Now, Construct 17−−√ on a number line, Step 1: Draw a number line with equal marks on either side. Step 2; Consider a point O at zero. Step 3: Mark a point A at 4 such that OA is 4. Step 4: Construct AB of unit length. Step 4: Join AB Step 5: Take OB as radius and intersect the number line at C. Step 6: Finally, C represents 17−−√ Note: Follow the steps carefully. If you miss any step in constructing the number on a number line, then you will get a wrong answer. A number line can extend infinitely in any direction. The left side of the number line is called the negative side and the right side is called the positive side.